OpenStudy (anonymous):

Inverse trigonometric functions(Calculus) u=Arctan x/a ans: du/dx=a/a^2+x^2 u= Arcsin x/a, a>0ans: du/dx= 1/sq. root a^2-x^2

OpenStudy (anonymous):

looks good to me

OpenStudy (anonymous):

we can go through it step by step if you like, as soon as the site stops having a fit

OpenStudy (anonymous):

yes pls satellite..

OpenStudy (anonymous):

i solved it but the answer is different from the the answer in the book

OpenStudy (anonymous):

first thing to know is that by the chain rule, if you want the derivative of \[f^{-1}(x)\] so find it by the following : \[\frac{d}{dx}f^{-1}(x)=\frac{1}{f'(f^{-1}(x))}\]

OpenStudy (anonymous):

that is because \[f(f^{-1}(x))=x\] and so \[\frac{d}{dx}f(f^{-1}(x))=1=f'(f^{-1}(x))\times \frac{d}{dx}f^{-1}(x)\] by the chain rule. solving this for \[\frac{d}{dx}f^{-1}(x)=\frac{1}{f'(f^{-1}(x))}\]\]g

OpenStudy (anonymous):

sorry site if freaking out on my tonight

OpenStudy (anonymous):

am satellite can i show my solution in the 2nd equation? and can you correct it pls..bec. i can understant better in application.

OpenStudy (anonymous):

sure go ahead.

OpenStudy (anonymous):